# Yuri Balashov

### v1

@@ 6.1 The Minimal Role of Instruction Tuning
After their initial pre-training LLMs normally undergo supervised instruction tuning which improves their responsiveness to specific requests. This is done on a mixture of mostly proprietary data, but there is little doubt that translation tasks are included in the mix. In any case, they are present in open-source instruction tuning datasets such as Aya (Singh et al., 2024). How important are they to the emergence of LLMs’ translation abilities?
@@
GPT-3 continued with another French sentence instead of translating (Reynolds and McDonell, 2021).
@@
Although there is still some dependence on the prompting details (Hendy et al., 2023; Peng et al., 2023; Sizov et al., 2024), in the latest LLMs suboptimal prompts tend to yield overly generous rather than incorrect responses to translation tasks. Some models need to be restrained with longer prompts in order to limit their output to what is needed (e.g. Balashov et al. 2025).
@@
So: while it may be tempting to attribute LLMs’ translation capabilities to explicit instruction tuning, it likely plays only a minor or supplementary role. Indeed, the translation capabilities of LLMs were already prominent in earlier, non-instruction-tuned models, such as GPT-2 and GPT-3. And they are stronger in non-instruction-tuned versions of more balanced multilingual models, such as Llama-3 and Llama-4, even if the latter are smaller in size. Due to their exposure to massive multilingual texts, LLMs inadvertently learn cross-lingual patterns and semantic mappings. Instruction tuning may enhance model responsiveness and specificity, but it does not fundamentally generate their cross-lingual capabilities. The end result is that, once prompted to translate, such a model can search its deeply learned space of multilingual representations and produce coherent translations, even for language pairs it has not been explicitly trained to translate in a supervised manner. Good prompts are more important than instruction tuning on translation examples.
@@ 6.2 Incidental Bilingualism in Training Data
In an earlier work, Eleftheria Briakou and colleagues (Briakou et al., 2023) approached this question by searching for “bilingual needles” in a haystack of PaLM’s (Chowdhery et al., 2022) training data. The authors hypothesized that incidental bilingualism—the unintentional inclusion of bilingual text within single training instances—is a key source of PaLM’s translation capabilities. In other words, the 540B-parameter model may have learned to translate from naturally occurring translations and mixed-language content in its massive training corpus (780 billion tokens of crawled text from web pages, social media, books, and so on), even though it was never given dedicated parallel data. The idea is that, while combing through its training corpus, PaLM may have stumbled upon numerous “needles in the haystack”: sentences and documents that appear side by side in different languages, essentially acting as implicit translation examples. These bilingual snippets were not explicitly labeled as “training examples” for translation; rather, they were incidental, arising naturally in the wild text. For example, a website might list a paragraph in English followed by the same paragraph in Spanish, or a forum post might include a quote in French with an English gloss. Briakou et al. suspected that such incidental bilingual data could be the primary source from which PaLM picked up its translation skills.
@@
Briakou and colleagues also observed variations on this theme: some documents used ISO language codes (e.g. ‘EN:’ or ‘FR:’) instead of full names, while others used language names in the native tongue (e.g., ‘Français’) to signal a French sentence). In some cases, the word ‘Translation’ in the target language was used as an indicator; e.g., a French text might begin with ‘Traduction:’ These are effectively natural prompts that inform a reader (or a model) that the following text is a translation of a preceding sentence.
@@
Exploiting these insights, Briakou et al. constructed prompts that mirror the patterns found in the training data and tested them on PaLM’s translation tasks. The results showed a marked improvement in PaLM’s zero-shot translation performance when using data-derived prompts. In particular, for translations out of English, using the newly discovered prompt formats (such as native language labels) boosted the average translation quality by about 14 chrF points compared to a generic prompting approach.999chrF (Popović, 2015) is a character-overlap $F$-score metric for translation quality; a 14-point gain is substantial. This indicates that PaLM had indeed learned to respond to those familiar bilingual cues: when the prompt matched patterns it saw during training, it produced significantly better translations. In practical terms, the model’s true translation ability was higher than initially thought. The “emergent” zero-shot performance was underestimated when users didn’t prompt the model in the right way. By prompting in a style consistent with the incidental training examples, PaLM could be coaxed to translate more accurately and consistently; for example, producing output in the correct target language more reliably. This highlights how context and presentation can govern LLM’s behavior: the relevant “knowledge” was latent in PaLM, yet only fully expressed when the trigger phrase matched its ingrained experience. It’s a bit like a person knowing a skill but needing the right cue or reminder to demonstrate it.
@@
There is, of course, no guarantee that the translation pairs mined from the internet are good. In fact, the general quality of the internet’s multilingual content is low (Kreutzer et al., 2022) and continues to degrade due to the rapidly increasing contamination from MT and LLM output (Thompson et al., 2024; Kocyigit et al., 2025). Briakou et al. performed an extrinsic evaluation of their makeshift corpus of 3.3M French-English sentence pairs extracted from PaLM’s translation instances by using them to train a standard NMT model from scratch and tested its performance. The resulting scores (37–38 BLEU) were not far from those (41 BLEU) for the same NMT trained on the full 40M-sentence WMT dataset. This means that PaLM’s training corpus implicitly contained a parallel corpus sufficient to train a decent translation system. The “needles” it swallowed were largely real and useful translation signals, not just noise. This reinforces the notion that PaLM had effectively learned from legitimate translation examples, so its translation ability is built on a foundation of authentic bilingual knowledge gleaned from the web.
@@
Another insight is how small a fraction of data can significantly influence behavior. Less than 1% of PaLM’s training instances were explicitly bilingual, yet those bits had an outsized impact on the model’s performance on translation tasks. This asymmetry points to the efficiency of pattern-learning in LLMs: when the model encounters a rare but highly informative pattern (like a sentence and its translation), it can leverage it to handle many other instances at inference time. It also means that omitted data could lead to omitted capabilities. If, hypothetically, PaLM’s (and, by extrapolation, other full-scale LLMs’) crawl of the web had missed all bilingual content, it might have shown far weaker translation abilities, at least until reaching enormous scale where other heuristics kick in. Thus, the structure and composition of training data directly shape the emergent “translation knowledge.”
@@ 6.3 Global and Local Learning
We have already noted (Section 5) that LLMs are capable of learning shared multilingual or language-neutral representations during training on semantically related monolingual documents, thanks to the awesome power of distributional semantics. In that respect, Global multilingual learning in LLMs is similar to cross-lingual representation alignment in mBERT and its variants, but on steroids. For example, real-world facts (dates, names, places) tend to appear repeatedly across languages, so the model continuously refines how these concepts are expressed in their respective monolingual contexts. If an LLM sees an article in English about climate change and another in German about ‘Klimawandel’ with overlapping terminology and structure, it can infer Global alignment between them. In general, when it encounters a text in language A that is very similar to a text in language B it saw earlier, the only way to reduce perplexity (prediction error) is to internally link those two contexts. Over many such occurrences, the model might cluster representations of sentences by meaning rather than by language. Importantly, this kind of Global learning (alignment, clustering) does not happen as a result of scanning single translation instances and adjusting the attention and other model weights in response. Rather, it happens in the course of repeated readjustment of the weights throughout multiple training steps.
@@
This kind of ongoing iteration between Local and Global learning is a natural and very helpful consequence of batch training which causes the model to update its weights after every training step. Once updated, the weights are used in the next iteration, continuously improving the model’s translation abilities (among other abilities it may learn in the process). Notably, the very same processes that are deployed in a much larger context window at inference are also operative in the multiple iterations in a narrower window during training.
The strategy of continuous iteration between “local” and “global” processes is deeply embedded in machine learning, optimization, and probabilistic modeling. This paradigm allows systems to gradually refine local parameters based on global structure or objectives, and vice versa, converging on a stable final state. The original Expectation-Maximization (EM) algorithm is a foundational framework for this interplay (Dempster et al., 1977). Many other techniques—from Latent Dirichlet Allocation (Blei et al., 2003) to statistical MT (Brown et al., 1993) to SentencePiece (Kudo and Richardson, 2018)—borrow this pattern.
So, on the balance of evidence and common wisdom, LLMs’ translation skills seem to emerge from a combination of these Local and Global learning processes. In practical terms, the model appears to leverage direct local correspondences when they are available, and fall back on more generalized global semantic alignment when they are not. This dual-sourced competence can explain why LLM translations sometimes resemble verbatim dictionary lookups and other times read like fluent paraphrases.
@@ 6.4.1 Style of Translation Outputs
If the model’s properties learned locally (memorized bilingual snippets) are driving a particular translation, the output should skew more literal and form-fixed, whereas globally learned cross-lingual representations will be more adaptive. For example, we expect LLMs to produce standard, idiomatic translations for common phrases that likely appeared in parallel form during training (e.g. proverb-like expressions, technical terms), but to generate more explicative or rephrased translations for novel or rare expressions.
Recent analyses of LLM outputs suggest that their translations tend to be literal for frequent source-target phrase pairs, yet more interpretative for inputs that were unlikely to have direct parallel examples. This aligns with manual analysis: in earlier studies, GPT models often rendered common idioms in a straightforward way but would explain obscure idioms in the target language when a ready equivalent is not known (a sign of on-the-fly alignment). For instance, Bureau Works’ analysis (Bureau Works, ) found that GPT-3 achieved 90% accuracy in translating Chinese idioms, outperforming traditional machine translation engines, which often produced literal and awkward translations. Similarly, Tang et al. (2024) demonstrated that GPT-4, when prompted appropriately, could generate high-quality, context-aware translations of East Asian idioms, surpassing commercial translation engines in both faithfulness and creativity.
These findings are further supported by comprehensive evaluations comparing GPT-4 to human translators. Yan et al. (2024) observed that GPT-4 tends toward overly literal translations and exhibits lexical inconsistency in some situations where human translators sometimes over-interpret context and introduce distinctly human “hallucinations” resulting from fatigue. This suggests that GPT-4’s translation behavior varies depending on the familiarity of the source phrases, being more literal with common expressions and more interpretative with less familiar ones.
@@ 6.4.2 Dependence on Model Scale and Data
Local vs. global learning should result in different scaling behaviors. A smaller LLM (with fewer parameters or less training data) is expected to rely more on memorized translation examples, while a larger model can compensate via broader semantic abstraction. Indeed, the ablation studies by Briakou et al. (2023) (Section 6.2 above) find that removing incidental parallel data causes a dramatic performance drop in a 1B-parameter model but a smaller drop in an 8B model. We predict this trend to extend to even bigger models: as we scale to tens or hundreds of billions of parameters, translation quality will increasingly survive the removal or absence of direct parallel pairs, because larger models should be able to infer better mappings through global context. Conversely, very small models might fail to translate at all without seeing explicit examples. Recent work on “English-only” pre-training corpora underscores this point: even tiny fractions of foreign text (well under 1% of tokens) greatly boost an English LLM’s cross-lingual performance, and target-language accuracy correlates strongly with how much of that language appeared in training. In other words, when a model lacks capacity for deep semantic abstraction, it needs those local signals; when it has sufficient capacity, it can achieve a form of translation by aligning semantics across languages (analogous to how multilingual BERT finds a shared embedding space). A concrete empirical test here is scaling ablation: train or fine-tune a series of models of increasing size on a fixed corpus with and without parallel examples, then measure translation fidelity. The dual-learning theory predicts an interaction: larger models should show relatively smaller gaps between the “with pairs” and “without pairs” conditions (relying on global learning to fill the gap), whereas smaller models will be crippled without explicit pairs. If instead all model sizes suffer proportionally or a large model is equally impaired by removal of parallel data, that would indicate a single-mode (or at least scale-invariant) learning process, contradicting the hypothesis.
@@ 6.4.3 Generalization to Unseen Language Pairs and Domains
Global semantic learning implies that an LLM can perform translation even between language pairs or domains it never saw aligned in training, by bridging via meaning. By contrast, purely local learning would struggle in truly novel transfers. We therefore predict that zero-shot translation—especially for unusual language combinations or highly specialized topics—is powered mostly by global mechanism. Empirically, this can be tested by looking at how well an LLM translates between languages that had no overlapping parallel data in its training set. If global learning is real, the model should still handle such pairs (perhaps by mapping each language to an English-centric semantic space, effectively pivoting internally; see Sections 4.3 and 5.3 above). Indeed, prior research on multilingual NMT showed that a single model can learn to bridge languages via an implicit “interlingua” when direct pairs are absent (Johnson et al., 2017). For LLMs, we have contemporary support: the latest GPT and Claude models reportedly perform decently on language pairs like Gujarati-Swahili or Bengali-Turkish (Enis and Hopkins, 2024) even though it is very unlikely they had parallel data for those pairs; they rely on internal representations and possibly English as a latent pivot.
@@
Another area to examine is the models’ capacity for domain and/or stylistic adaptation. An LLM fine-tuned on general web text might never see, say, legal contract sentences in both French and English in the same context. Yet, globally, it learns the legal terminology and style in each language. The dual-origin view predicts the model can translate legal text reasonably well (by aligning the semantics of legal expressions learned separately in French and English), but if a specific provision was present verbatim in a training bilingual document, the model might output the known translation verbatim. In evaluations, this would appear as high accuracy on common legal phrases (perhaps even outperforming human translators on consistency) but occasional weird literalness or errors on clauses that require more creative lexical adaptation (since global learning doesn’t provide a direct template). Testing on domain-specific parallel sets that were not in the training mix (to our best knowledge) could reveal this pattern. If observed, it would further support the idea that LLMs have both a “memory” for seen translations and a more dynamic ability to translate via understanding.
@@
In sum, treating LLMs’ translation abilities as having dual origins—part “memory-based,” part emergent alignment—provides a plausible explanation for the mixture of fluent generalization and oddly specific translations these models produce. It also aligns with what we know about scaling laws and training data: bigger models behave more “multilingually general,” and small ones cling to their few examples. The three predictions above offer avenues to verify this account. In the next section, I add some details on how one might design experiments to operationalize these predictions and more rigorously test the Local vs. Global learning hypothesis.

### v3

@@ 6.1 The Minimal Role of Instruction Tuning
After their initial pre-training LLMs normally undergo supervised instruction tuning which improves their responsiveness to specific requests. This is done on a mixture of mostly proprietary data, but there is little doubt that translation tasks are included in the mix. In any case, they are present in open-source instruction tuning datasets such as Aya (singh_aya_2024). How important are they to the emergence of LLMs’ translation abilities?
@@
GPT-3 continued with another French sentence instead of translating (reynolds_prompt_2021).
@@
Although there is still some dependence on the prompting details (hendy_how_2023; peng_towards_2023; sizov-etal-2024-analysing), in the latest LLMs suboptimal prompts tend to yield overly generous rather than incorrect responses to translation tasks. Some models need to be restrained with longer prompts in order to limit their output to what is needed (e.g. balashov-etal-2025-translation).
@@
So: while it may be tempting to attribute LLMs’ translation capabilities to explicit instruction tuning, it likely plays only a minor or supplementary role. Indeed, the translation capabilities of LLMs were already prominent in earlier, non-instruction-tuned models, such as GPT-2, GPT-3, and BLOOM. And they are stronger in non-instruction-tuned versions of more balanced multilingual models, such as Llama-3 and Llama-4, even if the latter are smaller in size. Due to their exposure to massive multilingual texts, LLMs inadvertently learn cross-lingual patterns and semantic mappings. Instruction tuning may enhance model responsiveness and specificity, but it does not fundamentally generate their cross-lingual capabilities. The end result is that, once prompted to translate, such a model can search its deeply learned space of multilingual representations and produce coherent translations, even for language pairs it has not been explicitly trained to translate in a supervised manner. Good prompts are more important than instruction tuning on translation examples.
@@ 6.2 Incidental Bilingualism in Training Data
In an earlier work, Eleftheria Briakou and colleagues (briakou_searching_2023) approached this question by searching for “bilingual needles” in a haystack of PaLM’s (chowdhery_palm_2022) training data. The authors hypothesized that incidental bilingualism—the unintentional inclusion of bilingual text within single training instances—is a key source of PaLM’s translation capabilities. In other words, the 540B-parameter model may have learned to translate from naturally occurring translations and mixed-language content in its massive training corpus (780 billion tokens of crawled text from web pages, social media, books, and so on), even though it was never given dedicated parallel data. The idea is that, while combing through its training corpus, PaLM may have stumbled upon numerous “needles in the haystack”: sentences and documents that appear side by side in different languages, essentially acting as implicit translation examples. These bilingual snippets were not explicitly labeled as “training examples” for translation; rather, they were incidental, arising naturally in the wild text. For example, a website might list a paragraph in English followed by the same paragraph in Spanish, or a forum post might include a quote in French with an English gloss. Briakou et al. suspected that such incidental bilingual data could be the primary source from which PaLM picked up its translation skills.
@@
Briakou and colleagues also observed variations on this theme: some documents used ISO language codes (e.g. ‘EN:’ or ‘FR:’) instead of full names, while others used language names in the native tongue (e.g., ‘Français’ to signal a French sentence). In some cases, the word ‘Translation’ in the target language was used as an indicator; e.g., a French text might begin with ‘Traduction:’. These are effectively natural prompts that inform a reader (or a model) that the following text is a translation of a preceding sentence.
@@
Exploiting these insights, Briakou et al. constructed prompts that mirror the patterns found in the training data and tested them on PaLM’s translation tasks. The results showed a marked improvement in PaLM’s zero-shot translation performance when using data-derived prompts. In particular, for translations out of English, using the newly discovered prompt formats (such as native language labels) boosted the average translation quality by about 14 chrF points compared to a generic prompting approach.888chrF (popovic_chrf_2015) is a character-overlap $F$-score metric for translation quality; a 14-point gain is substantial. This indicates that PaLM had indeed learned to respond to those familiar bilingual cues: when the prompt matched patterns it saw during training, it produced significantly better translations. In practical terms, the model’s true translation ability was higher than initially thought. The “emergent” zero-shot performance was underestimated when users didn’t prompt the model in the right way. By prompting in a style consistent with the incidental training examples, PaLM could be coaxed to translate more accurately and consistently; for example, producing output in the correct target language more reliably. This highlights how context and presentation can govern LLM’s behavior: the relevant “knowledge” was latent in PaLM, yet only fully expressed when the trigger phrase matched its ingrained experience. It’s a bit like a person knowing a skill but needing the right cue or reminder to demonstrate it.
@@
There is, of course, no guarantee that the translation pairs mined from the internet are good. In fact, the general quality of the internet’s multilingual content is low (kreutzer_quality_2022) and continues to degrade due to the rapidly increasing contamination from MT and LLM output (thompson_shocking_2024; kocyigit_overestimation_2025). Briakou et al. performed an extrinsic evaluation of their makeshift corpus of 3.3M French-English sentence pairs extracted from PaLM’s translation instances by using them to train a standard NMT model from scratch and tested its performance. The resulting scores (37–38 BLEU) were not far from those (41 BLEU) for the same NMT trained on the full 40M-sentence WMT dataset. This means that PaLM’s training corpus implicitly contained a parallel corpus sufficient to train a decent translation system. The “needles” it swallowed were largely real and useful translation signals, not just noise. This reinforces the notion that PaLM had effectively learned from legitimate translation examples, so its translation ability is built on a foundation of authentic bilingual knowledge gleaned from the web.
@@
Another key insight is that even a small fraction of data can exert a disproportionate influence on behavior. Less than 1% of PaLM’s training instances were explicitly bilingual, yet those bits had an outsized impact on the model’s performance on translation tasks. This asymmetry points to the efficiency of pattern-learning in LLMs: when the model encounters a rare but highly informative pattern (like a sentence and its translation), it can leverage it to handle many other instances at inference time. It also means that omitted data could lead to omitted capabilities. If, hypothetically, PaLM’s (and, by extrapolation, other full-scale LLMs’) crawl of the web had missed all bilingual content, it might have shown far weaker translation abilities, at least until reaching enormous scale where other heuristics kick in. Thus, the structure and composition of training data directly shape the emergent “translation knowledge.”
@@ 6.3 Global and Local Learning
We have already noted (Section 5) that LLMs are capable of learning shared multilingual or language-neutral representations during training on semantically related monolingual documents, thanks to the awesome power of distributional semantics. In that respect, Global multilingual learning in LLMs is similar to cross-lingual representation alignment in mBERT (devlin_bert_2019), mBART (liu-etal-2020-multilingual-denoising), and its variants, [^10] And, indeed, in earlier, pre-LLM static-embedding settings; e.g. mikolov2013exploitingsimilaritieslanguagesmachine exploited the potential of efficient learning of linear mappings between the monolingual embedding spaces in simple frameworks such as Skip-gram and CBOW. I am grateful to Michael Carl for reminding me of this important historical precedent. but on steroids. For example, real-world facts (dates, names, places) tend to appear repeatedly across languages, so the model continuously refines how these concepts are expressed in their respective monolingual contexts. If an LLM sees an article in English about climate change and another in German about ‘Klimawandel’ with overlapping terminology and structure, it can infer Global alignment between them. In general, when it encounters a text in language A that is very similar to a text in language B it saw earlier, the only way to reduce perplexity (prediction error) is to internally link those two contexts. Over many such occurrences, the model might cluster representations of sentences by meaning rather than by language. Importantly, this kind of Global learning (alignment, clustering) does not happen as a result of scanning single translation instances and adjusting the attention and other model weights in response. Rather, it happens in the course of repeated readjustment of the weights throughout multiple training steps.
@@
This kind of ongoing iteration between Local and Global learning is a natural and very helpful consequence of batch training, in which the model updates its weights after every step. In stochastic mini-batch training, each step’s gradient is computed over a shuffled mini-batch that typically mixes examples from different languages and domains. When a batch happens to contain both an occasional local bilingual cue and many globally related monolingual fragments, the single update aggregates these signals, simultaneously strengthening direct mappings and broader cross-lingual structure. Because mini-batches (often hundreds to thousands of sentences) may be reshuffled and resampled across successive steps and epochs, these heterogeneous mixtures can recur with different neighbors, so weights adjusted on one step are immediately put to work in the next one, closing the loop between Local and Global learning. Once updated, the weights are used in the next iteration, continuously improving the model’s translation abilities (among other abilities it may learn in the process). Notably, the same local processes are consistently employed across all stages of training.
In summary, then:
  • Local learning: acquisition from bilingual signals that co-occur within a single training context window (e.g., an English sentence followed shortly by its translation).
  • Global learning: alignment of semantically related monolingual content distributed across the training corpus, not necessarily co-occurring in a single window.
  • The two interact iteratively during batch training.
The strategy of continuous iteration between “local” and “global” processes is deeply embedded in machine learning, optimization, and probabilistic modeling. This paradigm allows systems to gradually refine local parameters based on global structure or objectives, and vice versa, converging on a stable final state. The original Expectation-Maximization (EM) algorithm is a foundational framework for this interplay (dempster_maximum_1977). Many other techniques—from Latent Dirichlet Allocation (blei_latent_2003) to statistical MT (brown_mathematics_1993) to SentencePiece (kudo_sentencepiece_2018)—borrow this pattern.
So, on the balance of evidence and common wisdom, LLMs’ translation skills seem to emerge from a combination of these Local and Global learning processes. Over the full course of training, the model appears to exploit direct local correspondences when available, and to rely on broader global semantic alignment when they are not. This dual-sourced competence can explain why LLM translations sometimes resemble verbatim dictionary lookups and other times read like fluent paraphrases.
@@ 6.4.1 Style of Translation Outputs
If the model’s properties learned locally (memorized bilingual snippets) are driving a particular translation, the output should skew more literal and form-fixed, whereas globally learned cross-lingual representations will be more adaptive. For instance, we expect LLMs to produce standard, idiomatic translations for common phrases that likely occurred in parallel form during training (e.g., proverb-like expressions or technical terms drawn from bilingual glossaries), but to generate more explanatory or rephrased translations for novel or rare expressions.
Recent analyses of LLM outputs suggest that their translations tend to be literal for frequent source-target phrase pairs, yet more interpretative for inputs that were unlikely to have direct parallel examples. This aligns with manual analysis: in earlier studies, GPT models often rendered common idioms in a straightforward way but would explicate obscure idioms in the target language when a ready equivalent is not known (a sign of on-the-fly alignment). For instance, Bureau Works’ analysis (bureau_works_we_nodate) found that GPT-3 achieved 90% accuracy in translating Chinese idioms, outperforming traditional machine translation engines, which often produced literal and awkward translations. Similarly, tang_creative_2024 demonstrated that GPT-4, when prompted appropriately, could generate high-quality, context-aware translations of East Asian idioms, surpassing commercial translation engines in both faithfulness and creativity.
These findings are further supported by comprehensive evaluations comparing GPT-4 to human translators. yan_benchmarking_2024 observed that GPT-4 tends toward overly literal translations and exhibits lexical inconsistency in some situations where human translators sometimes over-interpret context and introduce distinctly human “hallucinations” resulting from fatigue. This suggests that GPT-4’s translation behavior varies depending on the familiarity of the source phrases, being more literal with common expressions and more interpretative with less familiar ones.
@@ 6.4.2 Dependence on Model Scale and Data
Local vs. global learning should result in different scaling behaviors. A smaller LLM (with fewer parameters or less training data) is expected to rely more on memorized translation examples, while a larger model can compensate via broader semantic abstraction. Indeed, the ablation studies by briakou_searching_2023 (Section 6.2 above) find that removing incidental parallel data causes a dramatic performance drop in a 1B-parameter model but a smaller drop in an 8B model. We predict this trend to extend to even bigger models: as we scale to tens or hundreds of billions of parameters, translation quality will increasingly survive the removal or absence of direct parallel pairs, because larger models should be able to infer better mappings through global context. Conversely, very small models might fail to translate at all without seeing explicit examples. In other words, when a model lacks capacity for deep semantic abstraction, it needs local bilingual signals; when it has sufficient capacity, it can achieve a form of translation by aligning semantics across languages (analogous to how multilingual BERT finds a shared embedding space). A concrete empirical test here is scaling ablation: train or fine-tune a series of models of increasing size on a fixed corpus with and without parallel examples, then measure translation fidelity. The dual-learning theory predicts an interaction: larger models should show relatively smaller gaps between the “with pairs” and “without pairs” conditions (relying on global learning to fill the gap), whereas smaller models will be crippled without explicit pairs. If instead all model sizes suffer proportionally or a large model is equally impaired by removal of parallel data, that would indicate a single-mode (or at least scale-invariant) learning process, contradicting the hypothesis.
@@ 6.4.3 Generalization to Unseen Language Pairs and Domains
Global semantic learning implies that an LLM can perform translation even between language pairs or domains it never saw aligned in training, by bridging via meaning. By contrast, purely local learning would struggle in truly novel transfers. We therefore predict that zero-shot translation—especially for unusual language combinations or highly specialized topics—is powered mostly by global mechanism. Empirically, this can be tested by looking at how well an LLM translates between languages that had no overlapping parallel data in its training set. If global learning is real, the model should still handle such pairs (perhaps by mapping each language to an English-centric semantic space, effectively pivoting internally; see Sections 4.3 and 5.3 above). In fact, previous research on multilingual NMT showed that a single encoder-decoder model can learn to bridge languages through an implicit “interlingua” when direct pairs are absent from its training data (johnson_googles_2017). For LLMs, we have contemporary support: the latest GPT and Claude models reportedly perform decently on language pairs like Gujarati-Swahili or Bengali-Turkish (enis_llm_2024) even though it is very unlikely they had parallel data for those pairs; they rely on internal representations and possibly English as a latent pivot.
Recent experiments with toy transformer models trained on highly-structured knowledge graph-based multilingual synthetic factual datasets reported in blum2025rosettastoneunificationforces further suggest that larger models should better resist ablations of parallel data if training mixtures suppress language-as-shortcut signals (e.g., balanced attribute frequencies, tokenization that blurs script cues).
@@
Another area to examine is the models’ capacity for domain and/or stylistic adaptation. An LLM trained on general web text might never see, say, legal contract sentences in both French and English in the same context. Yet, globally, it learns the legal terminology and style in each language. The dual origin view predicts the model can translate legal text reasonably well (by aligning the semantics of legal expressions learned separately in French and English), but if a specific provision was present verbatim in a training bilingual document, the model might output the known translation verbatim. In evaluations, this would appear as high accuracy in common legal phrases (perhaps even outperforming human translators on consistency) but occasional weird literalness or errors on clauses that require more creative lexical adaptation (since global learning does not provide a direct template). Testing on domain-specific parallel sets that were not in the training mix (to our best knowledge) could reveal this pattern. If observed, it would further support the idea that LLMs have both a “memory” for seen translations and a more dynamic ability to translate via understanding.
@@
In sum, treating LLMs’ translation abilities as having dual origins—part “memory-based,” part emergent alignment—provides a plausible explanation for the mixture of fluent generalization and oddly specific translations these models produce. It also aligns with what we know about scaling laws and training data: bigger models behave more “multilingually general,” and small ones cling to their few examples. The three predictions above offer avenues to verify this account. In the next section, I add some details on how one might design experiments to operationalize these predictions and more rigorously test the dual origin hypothesis.